OpenStudy (anonymous):

x^2m(x^2m-2x^m+3) = ? Help..!

OpenStudy (anonymous):

with explanation. :)

OpenStudy (anonymous):

do u just wnt to simplify the expression ?

OpenStudy (anonymous):

@Rahammii

OpenStudy (unklerhaukus):

Is this your question?\[x^{2m}(x^{2m-2}x^{m+3}) = ? \]

OpenStudy (anonymous):

I think the question is\[x^{2m}(x^{2m}-2x^m+3)=?\]

OpenStudy (anonymous):

yup that is the question ^_^ mukushla

OpenStudy (anonymous):

what is the answer?

OpenStudy (marinos):

Do you need to simplify or factorize the expression ? Or is it an equation to be solved, in which case I assume you need it to be equal to zero ?

OpenStudy (anonymous):

simplify :)

OpenStudy (unklerhaukus):

\[a(b+c+d)=a(b)+a(c)+a(d)\] \[x^{2m}(x^{2m}-2x^m+3)=x^{2m}\times (x^{2m})+x^{2m}\times(-2x^m)+x^{2m}\times (3)=?\]